# LayerShuffle: Enhancing Robustness in Vision Transformers by Randomizing Layer Execution Order

### v1

@@ 4.2 Removing layers during test time
For arbitrary execution (Figure 1(c)), LayerShuffle is the only approach that succeeds, with the average accuracy improving as the number of layers is increased. LayerDrop does not perform well regardless of the number of layers in the model. A noteworthy detail is the comparable high average accuracy of the fully retrained baseline with 3 layers. Given the low performance of the refined models with 6 and 9 layers, as well as that there are only 6 possible permutations for 3 layers, the most likely explanation is that one of the 5 random permutations evaluated for the model was the original layer execution order the model has been trained for, i.e.  [1,2,3] therefore skewing the achieved accuracy in this case.
In conclusion, we find that our proposed approach has similar test-time scaling capabilities as LayerDrop, while still ensuring robustness towards arbitrary layer execution orders.
@@ 4.3 Analysis of intermediate network representations
To gain a deeper insight into how information is encoded in our trained models, we compute Uniform Manfold Approximation and Projection (UMAP) (McInnes et al., 2018) embeddings of the outputs of a particular attention module, where we color-code all output vectors based on the position the module held in the network when producing this output.
Figure 4: UMAP embedding of the output of a layer trained with shuffling execution order. The layer produces outputs in different subspaces of the latent space depending on their current position in the network. Lighter colors indicate layer positions closer to the input; layer positions close to the output are shown in dark colors. Note that information about the layer position has not been presented to the UMAP algorithm and the visual proximity of points with similar layer position is an emergent feature of our training approaches.
@@
For the UMAP analysis, we present 1,000 randomly sampled images from the ImageNet2012 validation set to a LayerShuffle-trained model. While we use an evaluation batch size of 1 image and record all outputs of a single, previously selected layer, we randomly permute the execution order of layers such that the selected layer changes position in the network during every forward path. After the layer output vectors for all 1,000 images have been recorded, a UMAP reduction of the output space to 2D is performed.
@@
Figure 4 shows the two-dimensional UMAP-embeddings of the output space of one particular layer in a trained LayerShuffle transformer. The current position of the layer in the network when producing a given output is color-coded from light position close to the input to dark (position close to the output). Note that this information about the layer position has not been presented to the UMAP algorithm. The plot shows that, while there is no sharp separation between outputs generated at different positions in the network, the layer clearly adapts to its current position and extracts different features for different positions in the network. A further interesting observation is the very distinct collection of points for layer positions close to the input, which are detached from the remaining manifold of points, which suggests that this initial stage, where early low-level features are extracted, requires special treatment of the layer.
4.4 Merging models with arbitrary execution order
Being robust against permuting the layer execution order, opens interesting other possibilities such as model merging, i.e. creating a new model from the layers of several identically trained models. The underlying rationale is that such merged models could also occur in a distributed setting, where compute nodes, whose layers have been trained as part of distinct models, but with the same training process, could form ad-hoc models together.
To construct merged models, where each layer stems from a different model, we require $12$ models for the $12$ layers of the ViT-B/16. We therefore train 7 more networks for our LayerShuffle approach and the baseline. Subsequently, we create 100 merged models (out of $12!$ possible combinations) by randomly sampling from these models for our proposed approach as well as the baseline respectively (models are not mixed between approaches). As mentioned previously, layers are sampled in such a way that no two layers in a merged model stem from the same model. We then evaluate the validation accuracy of all 100 models for both approaches.
Table 5 summarises the results. The merged baseline model ViT-B/16 deteriorates from $82.61\%$ average accuracy to $1.87\%$ (despite sequential layer execution as required by the model) making the resulting merged model effectively unusable. The merged LayerShuffle models, on the other hand, perform slightly below the original model with an average accuracy of $59.68\%$ as opposed to the $62.77\%$ of the latter. Less surprisingly, merged models show a higher standard deviation at $1.15$ percentage points for the merged models vs. $0.41$ percentage points for the original ones as merged models do not contain any two layers that have been trained together, which makes their performance vary more. We can further improve performance by ensembling the 12 models trained with LayerShuffle, using the average of their output logit vectors. Such neural network ensembles often reach a better performance (Hansen & Salamon, 1990), which is also the case here with a significant improvement and an accuracy of $69.20\%$ for LayerShuffle Ensemble.
Visualizing the accuracies of all achieved models in a swarm plot paints a similar picture (Figure 5). While most merged LayerShuffle achieve an accuracy close to the average accuracy of the original models, for the baseline, all merged models exhibit very low accuracies. In conclusion, we find that permuting layer order during training enables the construction of merged (or ”Frankenstein”) vision transformers, where each layer of the transformer can be taken from a different model, as long as all models have been refined from the same base model on the same data.
Table 2: Accuracy of merged ViTs. Merged LayerShuffle achieve an accuracy close to the average accuracy of the original models, while for the baseline, the merged model exhibits very low accuracy. Ensembles of LayerShuffle models show clear improvement over single models.
Model Top-1 Acc. Top-5 Acc. layer order
ViT-B/16 merged $1.87\pm 6.51\%$ $4.53\pm 11.82\%$ sequential
LayerShuffle merged $59.68\pm 1.15\%$ $82.16\pm 1.03\%$ arbitrary
LayerShuffle Ensemble $69.20\%$ $88.93\%$ arbitrary
Figure 5: ImageNet2012 validation accuracies of merged models for LayerShuffle and ViT-B/16 models. Enforcing random layer execution order during training permits the construction of ViT Frankenstein models. The performance of merged models composed of Baseline models on the other hand breaks down.
@@ 5 Discussion and Future Work
This paper presented a new approach called LayerShuffle, which enabled vision transformers to be robust to arbitrary order execution, pruning at test time, as well as adhoc-construction of merged models. For sequential execution, LayerShuffle performs on average only slightly worse than the LayerDrop approach but is the only method that works when the layer execution is arbitrary. A latent space analysis confirmed that layers of models trained with LayerShuffle adjust their output depending on which position they hold in the network. Finally, we investigated whether it is possible to build merged models from the models trained with LayerShuffle and found the performance of the built merged models to be only slightly less than the performance of our trained models, contrary to the baseline, where virtually all merged models delivered very poor performance.
That LayerShuffle works well is surprising and here we have only scratched the surface of its effect on the trained neural networks. To gain a deeper understanding, further analysis will include an inspection of the norms of the outputs of multi-head-attention and multi-layer-perceptron layers. This type of investigation could reveal whether layers learn to switch off their output for inputs they cannot process such that further relevant processing can occur in a suitable downstream layer to which the data is relayed through the attention module’s residual connections. Inspecting the model’s attention maps and embedding the intermediate latent vectors of all layers at all positions in a single two-dimensional embedding might yield further insights as well.
In the future, these properties could make LayerShuffle-trained models ideal candidates to be distributed over a number of very loosely coupled compute nodes to share the computational load of model inference. Given the enormous engineering, financial and logistical effort as well as the environmental impact (Strubell et al., 2020) of building and maintaining datacenters for state-of-the-art deep learning approaches on the one hand, as well as the large amount of available, but scattered compute through existing smartphones, laptop computers, smart appliances and other edge devices on the other hand, approaches that allow ad-hoc neural networks performing inference together could be of great impact. We therefore consider the deployment and orchestration of our trained models onto an actual set of edge devices and the practical implementation of the inference process on a network of such devices, likely by combining our approach with previously proposed frameworks to address this issue (Gacoin et al., 2019), a very promising direction of future research.
@@ Acknowledgements
We thank the members of the REAL and Creative AI group for useful discussions. This work was supported by a research grant (40575) from VILLUM FONDEN.

### v2

@@ 4.2 Removing layers during test time
For arbitrary execution (Figure 1(c)), LayerShuffle is the only approach that succeeds, with the average accuracy improving as the number of layers is increased. LayerDrop does not perform well regardless of the number of layers in the model.
A noteworthy detail is the comparable high average accuracy of the fully retrained baseline with 3 layers. Given the low performance of the refined models with 6 and 9 layers, as well as that there are only 6 possible permutations for 3 layers, the most likely explanation is that one of the 5 random permutations evaluated for the model was the original layer execution order the model has been trained for, i.e.  [1,2,3] therefore skewing the achieved accuracy in this case. In conclusion, we find that our proposed approach has similar test-time scaling capabilities as LayerDrop, while still ensuring robustness towards arbitrary layer execution orders.
@@ 4.3 Analysis of intermediate network representations
(a) Baseline
(b) Refined w. LayerShuffle
(c) Baseline
(d) Refined w. LayerShuffle
Figure 4: UMAP-projected embeddings and contributions to model prediction (estimated distribution of normalized L2 norms of class token) of layer outputs trained with shuffling execution order, baseline for comparison. Contrary to the baseline (a), the layer for a LayerShuffle-trained network (b) produces outputs in different subspaces of the latent space depending on their current position in the network. Darker colors indicate layer positions closer to the input; layer positions close to the output are shown in light colors. While layers in the baseline model overall contribute equally to the predictive output of the model, regardless of their current position in the network (c), the contribution of layers in the LayerShuffle-trained model’s prediction (d) varies based on the distance to it’s original position in the networks. Refinement of the model conditions its layers to only contribute to the overall predictive output if the received input lies within the layers learned distributions of inputs.
@@
To gain a deeper insight into how information is encoded in the models, we conduct two experiments. First, we compute Uniform Manifold Approximation and Projection (UMAP) (McInnes et al., 2018) embeddings of the entire output of a particular attention module (i.e. combined self-attention and feed-forward layers), where we color-code all output vectors based on the position the module held in the network when producing this output. In more detail, we concatenate all patch tokens of a single image together with the class token as a single vector, and use this representation as a single state vector in our compression. To extract a sufficient number of these state vectors, we present 1,000 randomly sampled images from the ImageNet2012 validation set to a LayerShuffle-trained model. While we use an evaluation batch size of 1 image and record all outputs of a single, previously selected layer, we randomly permute the execution order of layers such that the selected layer changes position in the network during every forward path. After the layer’s output vectors for all 1,000 images have been recorded, a UMAP reduction of the output space to 2D is performed.
Second, in order to investigate how much each layer contributes to the final classifier output when deployed in different positions within the model, we compute the L2-Norm of the class-token of each layer output. We correct for the contribution of previous layers by subtracting the class token of the previous layer before computing the token’s norm. That way, we consider solely the additive contribution of the layer to the class prediction of the model. We collect these token norms for all network layers as we shuffle their position while presenting 1000 randomly sampled input images in an identical manner as in the previous experiment.
@@
Finally, to establish a baseline, we extract both representations for the original ViT-B/16 weights Dosovitskiy et al. (2020) as well. Figure 4(b) shows the obtained visualizations for the original ViT-B/16 model acting as a baseline as well as our model refined with LayerShuffle. In more detail, Figures 4(a) and 4(b) show the UMAP embeddings of a single layer’s output for both the baseline and our model. The current position of the layer in the network when producing a given output is color-coded from dark (position close to the input) to light (position close to the output). Note that this information about the layer position has not been presented to the UMAP algorithm. Apart from rough trends, no clear ordering of the space is visible for the baseline (Figure 4(a)). For LayerShuffle, while there is no sharp separation between outputs generated at different positions in the network, the layer clearly adapts to its current position and extracts different features for different positions in the network (Figure 4(b)).
A further interesting observation is the very distinct collection of points for layer positions close to the input, which are detached from the remaining manifold of points. This results suggests that extracting low-level features, requires special treatment.
Figures 4(c) and 4(d) show the distributions on the normalized L2-norm of additive contributions to class tokens for different layer positions in the transformer for both the baseline and our model. Each x-axis in the plot corresponds to a single layer of the network, where position of the layer in the network is color-coded again from close to the input (dark) to close to the output (light). x-axes are also ordered corresponding to the layer’s original position in the pre-trained model, where the order of layers is top to bottom. We can see that for the baseline model norms are basically spread out over the whole range. This implies that layers in the baseline model overall contribute evenly to the predictive output of the model, regardless of their current position in the network.
On the other hand, the ridge plot gathered from layer outputs of the model refined with our method paints a different picture. The norm of attention modules output and therefore it’s contribution to the model’s prediction varies based on the distance to it’s original position in the networks. Modules which were originally closer to the input (x-axes on top of the plot) often show larger contributions to the predictive output of the model when on positioned closer to the input and vice versa. This indicates that our refinement of the model conditions its layers to contribute to the overall predictive output if the received input lies within the layers learned distributions of inputs (i.e. the layer is close at a position assigned to it in the original pre-trained network), and withhold or reduce their output otherwise. This is also in line with recent work conducted in parallel Sun et al. (2024), which frames transformer layers as incrementally refining a rough sketch of the model’s output, an iterative process which is enabled by the transformer’s extensive utilization of skip-connections.
In conclusion, our analysis indicates that refining networks with LayerShuffle makes vision transformers robust to arbitrary execution orders as it trains the layers to solely add to the models contribution if the layer input is in-distribution and reduce their output otherwise, in which case the model’s skip-connection forwards the out-of-distribution output to the subsequent layer.
@@ 5 Discussion and Future Work
This paper presented a new approach called LayerShuffle, which enabled vision transformers to be robust to arbitrary order execution as well as pruning at test time. For sequential execution, LayerShuffle performs on average only slightly worse than the LayerDrop approach but is the only method that works when the layer execution is arbitrary.
Our analysis confirmed that layers of models trained with LayerShuffle adjust their output depending on which position they hold in the network. Furthermore, our results indictate that refining networks with LayerShuffle trains the layers to only contribute to the model’s class prediction if the layer input is in-distribution and reduce their output otherwise, in which case the layer’s skip-connection forwards the barely modified out-of-distribution embedding to the subsequent layer.
In the future, these properties could make LayerShuffle-trained models ideal candidates to be distributed over a number of very loosely coupled compute nodes to share the computational load of model inference. Given the enormous engineering, financial and logistical effort as well as the environmental impact (Strubell et al., 2020) of building and maintaining datacenters for state-of-the-art deep learning approaches on the one hand, as well as the large amount of available, but scattered compute through existing smartphones, laptop computers, smart appliances and other edge devices on the other hand, approaches that allow distributed neural networks to perform inference could be of great impact. We therefore consider the deployment and orchestration of our trained models onto an actual set of edge devices and the practical implementation of the inference process on a network of such devices, likely by combining our approach with previously proposed frameworks to address this issue (Gacoin et al., 2019), a very promising direction of future research.
@@ Acknowledgements
We would like to thank Prof. Christian Igel for his insightful feedback on the manuscript. Furthermore we thank the members of the REAL and Creative AI lab for fruitful discussions. This work was supported by a research grant (40575) from VILLUM FONDEN.